# Sublattice-selective percolation on bipartite planar lattices

### v1

@@ III.2 Critical Exponents
In addition to the percolation threshold, we estimated the critical exponents for sublattice-selective percolation on the square lattice using appropriate finite-size scaling analysis Stauffer and Aharony (2018). We first consider the critical exponent $\nu$, which can be best estimated from the finite-size scaling of the gradient $R^{\prime}_{L}=dR_{L}/dp_{A}$ of $R_{L}(p_{A})$ near the percolation threshold,
$R^{\prime}_{L}\propto L^{1/\nu},$ (9)
We approximate $R^{\prime}_{L}$ from the slope of a linear regression within a $p_{A}$-range of $10^{-3}$ around the estimated value of $p_{A,c}$. The thus obtained estimates for $\nu$ are given in Tab. 1, and Fig. 7 shows bests fits to the numerical data for different values of $p_{B}$.
@@
The values that we obtain for $\nu$ scatter with no systematic dependence of $p_{B}$, and taking their average gives $\nu=1.3352(11)$, which is in good agreement with $\nu=4/3$ for the $d=2$ universality class. We also find no indication that the critical exponents $\beta$ and $\gamma$ differ from their conventional values for $d=2$. In order to access these exponents, we consider the finite-size scaling forms
@@
for the mean number $S$ of finite clusters. From appropriate data-collapse plots, one can adjust the values of $\beta$ and $\gamma$, based on our previous estimates for $\nu$ and $p_{A,c}$, in order to obtain the best collapse within the critical region. Using such an analysis, we find that the best fit values are in fact in accord with the known values for conventional percolation in $d=2$. This is illustrated for several values of $p_{B}$ in Fig. 8. We thus conclude from our finite-size analysis, that the critical exponents for sublattice-selective percolation are in agreement with the universal values for $d=2$ percolation, supporting previous findings Scholl and Binder (1980); Idogaki and Uryû (1982).
@@ IV Lieb lattice
As a further example, we examine sublattice-selective percolation on the Lieb lattice, for which the case of uniform site percolation ($p_{A}=p_{B}$) has recently has been considered Oliveira et al. (2021). Since the Lieb lattice is in fact a decorated square lattice, sublattice-selective percolation is in this case equivalent to mixed bond-site percolation on the square lattice, with $p_{A}$ ($p_{B}$) specifying the bond (site) occupation probability. In particular, in the limit $p_{A}=1$ ($p_{B}=1)$, we obtain conventional pure site (bond) percolation on the square lattice, and thus $p_{B,c}=p_{c}^{\mathrm{sq}}$ ($p_{A,c}=1/2$) along these lines, respectively (here, we used the fact that the percolation threshold $1/2$ for bond percolation on the square lattice has been exactly determined Kesten (1980)). This shows explicitly that due to the inequivalence of the two sublattices of the Lieb lattice, the phase diagram is not symmetric in the $(p_{A},p_{B})$-plane. In Ref. Tarasevich and Van der Mark (1999), the general case of mixed bond-site percolation was considered for various lattices, including the square lattice. The phase diagram was numerically estimated, but only to comparably low accuracy, so that here we employed the bipartite Newman-Ziff algorithm in order to obtain more accurate values for the threshold line for sublattice-selective percolation on the Lieb lattice. The numerical results are provided in Tab. 2, and the corresponding phase diagram is shown in Fig. 9. Comparing the value of $p_{A,c}$ obtained for $p_{B}=1$ to the exact value ($1/2$) confirms the accuracy of our results. For estimating these numbers, $C$ was chosen identical to the square lattice value from the previous section, resulting in similarly fast convergence. Additionally, the previous estimate Oliveira et al. (2021) for the conventional ($p_{A}=p_{B}$) site percolation threshold $p_{c}^{\mathrm{Lieb}}$ on the Lieb lattice was improved using the Newman–Ziff algorithm. Along this line, we find that $C=0.68$ yields better convergence (however, even then the exponent $X\approx 1/\nu$ is significantly lower). The obtained value $p_{c}^{\mathrm{Lieb}}=0.7397120(17)$ is consistent with the previous best estimate Oliveira et al. (2021).
@@ V.1 Percolation on a chain ($z=2$)
In this case, percolation occurs only at the singular point $(p_{A},p_{B})=(1,1)$. In order to extract the critical exponents upon approaching the percolation point along different lines within the $(p_{A},p_{B})$-plane, we next examine appropriate state functions.
A cluster containing $s$ occupied sites with one empty site to either side is called a $s$–cluster. The number of $s$–clusters per lattice site $n_{s}$ is a useful state function. Another way of looking at $n_{s}$ is that it is the probability of an arbitrary site to be the left-end site of an $s$–cluster. This is the case if the site to the left is empty, then $s$ consecutive sites are occupied, followed by another empty site. For $p_{A}=p_{B}=p$ the state function would be $n_{s}=(1-p)p^{2}(1-p)$, but for $p_{A}\neq p_{B}$ one has to consider whether the left end site is from the $A$ or $B$ sublattice, and whether $s$ is even or odd. The probability that an arbitrary site is from the $A$ or $B$ sublattice is equal to $1/2$, and from there on it is a matter of counting the $A$– and $B$–sublattice sites in the cluster to find
@@
for even $s$, where $\bar{p}_{A}=1-p_{A}$, $\bar{p}_{B}=1-p_{B}$. The probability that an arbitrary site is part of a $s$–cluster is $n_{s}\cdot s$, so the probability that an arbitrary site is part of any cluster is given by
@@
which approaches 1 for $(p_{A},p_{B})\rightarrow(0,0)$ and diverges upon approaching $(1,1)$. We obtain the critical exponent $\gamma$ from analyzing the singular behaviour of $S$ near $(1,1)$. Fixing $p_{B}=1$, the asymptotic singularity is obtained as
@@
In order to obtain the correlation length $\xi$, we first calculate the correlation function (pair connectivity function) $g(r)$, i.e., the probability that an arbitrary occupied site is in the same cluster as another site a distance $r$ apart. Thus, there have to be $r-1$ consecutive occupied sites in between. For $p_{A}=p_{B}=p$ this is just $2p^{r}$ (2 for both directions), but for $p_{A}\neq p_{B}$ the parity of $r$ and the starting site have to be accounted for. The starting site is on the $A$ sublattice with probability $\frac{p_{A}}{p_{A}+p_{B}}$ and on the $B$ sublattice with probability $\frac{p_{B}}{p_{A}+p_{B}}$. Accounting for the parity of $r$ and the special case $r=0$, we obtain
$g(r)=\begin{dcases}1,&r=0\\ \frac{4}{p_{A}+p_{B}}p_{A}^{(r+1)/2}p_{B}^{(r+1)/2},&\text{$r$ odd}\\ 2p_{A}^{r/2}p_{B}^{r/2},&\text{$r$ even}\end{dcases}.$ (17)
@@
The numerator in Eq. (18) can also be calculated and for $p_{B}=1$ scales as $(1-p_{A})^{-3}$ near the percolation threshold. Since $1/S$ scales as $(1-p_{A})$, the singularity of $\xi$ is thus obtained as
@@ V.2 Percolation threshold line for general $z$
The percolation condition on the Bethe lattice is that a chain of occupied sites reaches out to infinity. In other words, the system percolates if from some occupied site there exists a path of connected occupied sites that never ends (note that on the Bethe lattice such the path cannot form loops). If a chain of sites from a specific site up to some other site are occupied, at least one of the two neighbours going outwards also has to be occupied to have a percolating cluster. Hence, the mean number of occupied sites going outwards has to be at least one in order to form a percolating cluster. Furthermore, on the Bethe lattice the $A$ and $B$ sublattice alternate. The mean number of occupied sites going outwards can thus be calculated for $z$ neighbours by using two nested binomial distributions, one accounting for the mean number of $A$-sublattice sites and the other for the mean number of $B$-sublattice sites. This leads to the condition
@@
which can be simplified to yield the condition for the percolation transition
@@
Note that this result is also correct for $z=2$, treated in the previous section. The percolation threshold line for $z=3$ is at $p_{A}\>p_{B}=1/4$, which is plotted in Fig. 10. The larger $z$, the more area of the phase diagram is percolating, but here only $z=3$ will be further examined.
@@ V.3 Scaling exponents for $z=3$
To obtain $P$, we first calculate the probability of a path not leading to infinity. Let $w_{A}$ ($w_{B}$) be the probability that an arbitrary $A$($B$) sublattice site does not lead to infinity, respectively. An $A$ sublattice site does not lead to infinity if the site itself is not occupied (given by the probability $1-p_{A}$) or if the two neighbours leading outwards do not lead to infinity (given by the probability $p_{A}w_{B}^{2}$), and similarly for a $B$ sublattice site. Altogether, this gives the recursive expressions
@@
This nonlinear system of equations has polynomials of degree four as solutions. The trivial solution is $w_{A}=w_{B}=1$, while the other real solution, denoted by $w^{P}_{A},w^{P}_{B}$ is rather lengthy, and given explicitly as a function of $p_{A}$ and $p_{B}$ in the appendix A. We note that for larger coordination number $z$, the corresponding recursion equations lead to higher-order equations for $w_{A}$ and $w_{B}$, for which closed expressions for the roots are known to not exist from Galois theory. For $p_{A}\>p_{B}<1/4$ there exists no infinite cluster and thus the trivial solution applies, while for $p_{A}\>p_{B}\geq 1/4$ the other real solution is valid, i.e.,
$w_{A/B}=\begin{dcases}1,&4p_{A}\>p_{B}<1\\ w^{P}_{A/B}(p_{A},p_{B}),&4p_{A}\>p_{B}\geq 1\end{dcases}.$ (25)
@@
since the probability that the center site is from the $A$ or $B$ sublattice is $1/2$ and the probability that the site is occupied and at least one neighbour leads to infinity equals $p_{A/B}(1-w_{B/A}^{3})$. The strength of the percolating cluster is plotted in Fig. 11.
@@
In Fig. 12 cuts of $P$ at three different values of $p_{B}$ are shown, along with the function $26p_{A}p_{B}-3/$ 2, which yields the leading asymptotic of $P$ for conventional site-percolation on the Bethe lattice Stauffer and Aharony (2018). We thus find that near the percolation threshold line, i.e., for $p_{A}\gtrsim 1/(4p_{B}),p_{B}\geq 1/4$, the scaling behavior
@@
Next, the mean cluster size $S$ will be calculated. For this an arbitrary occupied starting site is considered and the mean mass of the cluster it belongs to is calculated. This approach is only valid below the percolation threshold line, i.e., for $14p_{A}p_{B}<$ 1, because otherwise the mass of the percolating cluster has to be accounted for as well. If the starting site belongs to the $A$($B$) sublattice, then let $S_{B}$ ($S_{A}$) be the contribution to the mean cluster size from each neighbour. The mean cluster size then is composed of the contribution from the center site plus the contribution from the three neighbours, so that
@@
The contributions $S_{A}$ and $S_{B}$ can be accessed recursively: If the starting site belongs to the $A$ sublattice, then a given neighbour only contributes a non–zero $S_{B}$ with probability $p_{B}$. The size of this contribution is the site itself plus $S_{A}$ for the two neighbours going outwards. With a similar reasoning for the $B$ sublattice, we thus obtain
@@
Both $S_{A}$ and $S_{B}$ diverge upon approaching the threshold condition $p_{A}p_{B}=\frac{1}{4}$. Hence, the behaviour of $S$ near the percolation threshold line can be described by
@@
The correlation function $g(r)=g(\sqrt{r_{t}})$, for $14p_{A}p_{B}<$ 1, gives the mean number of sites connected to a given occupied site at a topological distance $r_{t}$, corresponding to the Euclidean distance $r$. At $r_{t}=0$ the center site is already occupied, while at $r_{t}=1$ the site has 3 neighbours and from there on the number of neighbours doubles in each step, and the probabilities are otherwise the same as for the case of the chain, so that we obtain
@@ VI Conclusions
We presented an adapted Newman–Ziff algorithm to sublattice-selective percolation on bipartite lattices and applied it to accurately determine the percolation threshold lines for the square lattice and the planar Lieb lattice. The latter case relates to mixed bond-site percolation on the square lattice and our numerical results for the percolation threshold line refine previous estimates. Our numerical estimates for the critical exponents at the percolation transition are consistent with the universality class for percolation in $d=2$.
In addition, we examined sublattice-selective percolation on the Bethe lattice, which was in fact the geometry considered in the pioneering work on percolation by Flory in the context of polymerization Flory (1941). For this case, we obtained the exact percolation threshold line and critical exponents for $z=2$ and $z=3$, which are again in accord with the values for conventional percolation. Given that the sites of the square lattice have a uniform coordination number of $z=4$, it may be interesting to compare its phase diagram to the one of the Bethe lattice with $z=4$. Such a comparison is shown in Fig. 13, illustrating the substantially larger percolating regime for the latter case. Indeed, the absence of closed loops for a chain of connected sites leads to an overall prior appearance of the percolating cluster upon increasing the site occupation probabilities on the Bethe lattice. We checked explicitly, that a rescaling of the analytic expression for the percolation threshold line for the $z=4$ Bethe lattice to match the ends points at $p_{A}=1$ and $p_{B}=1$ does not fit the full numerical data.
Related to the last point, we note that the authors of Ref. Tarasevich and Van der Mark (1999) considered various proposed analytic expressions for the general functional form of the percolation threshold line for mixed bond-site percolation. They observed clear systematic deviations from the numerical data for low-coordinated lattices. Based on our refined data on the Lieb lattice, we can also exclude the general validity of these expressions for the square lattice. It thus remains an interesting open question for further research to derive analytical expressions for the percolation threshold line for sublattice-selective as well as mixed bond-site percolation, respectively. We hope that our refined numerical data will be valuable as a benchmark for such investigations.

### v2

@@ III.2 Critical exponents
In addition to the percolation threshold, we estimated the critical exponents for sublattice-selective percolation on the square lattice using appropriate finite-size scaling analysis Stauffer and Aharony (2018). We first consider the critical exponent $\nu$, which can best be estimated from the finite-size scaling of the gradient $R^{\prime}_{L}=dR_{L}/dp_{A}$ of $R_{L}(p_{A})$ near the percolation threshold,
$R^{\prime}_{L}\propto L^{1/\nu}.$ (9)
We approximate $R^{\prime}_{L}$ from the slope of a linear regression within a $p_{A}$ range of $10^{-3}$ around the estimated value of $p_{A,c}$. The thus obtained estimates for $\nu$ are given in Table 1, and Fig. 7 shows bests fits to the numerical data for different values of $p_{B}$.
@@
The values that we obtain for $\nu$ scatter with no systematic dependence of $p_{B}$, and taking their average gives $\nu=1.3352(11)$, which is in good agreement with $\nu=4/3$ for the $d=2$ universality class. We also find no indication that the critical exponents $\beta$ and $\gamma$ differ from their conventional values for $d=2$. To access these exponents, we consider the finite-size scaling forms
@@
for the mean number $S$ of finite clusters. From appropriate data-collapse plots, one can adjust the values of $\beta$ and $\gamma$, based on our previous estimates for $\nu$ and $p_{A,c}$, in order to obtain the best collapse within the critical region. Using such an analysis, we find that the best-fit values are in fact in accord with the known values for conventional percolation in $d=2$. This is illustrated for several values of $p_{B}$ in Fig. 8. We thus conclude from our finite-size analysis, that the critical exponents for sublattice-selective percolation are in agreement with the universal values for $d=2$ percolation, supporting previous findings Scholl and Binder (1980); Idogaki and Uryû (1982).
@@ IV Lieb lattice
As a further example, we examine sublattice-selective percolation on the Lieb lattice, for which the case of uniform site percolation ($p_{A}=p_{B}$) has recently been considered Oliveira et al. (2021). Since the Lieb lattice is in fact a decorated square lattice, sublattice-selective percolation is equivalent in this case to mixed bond-site percolation on the square lattice, with $p_{A}$ ($p_{B}$) specifying the bond (site) occupation probability. In particular, in the limit $p_{A}=1$ ($p_{B}=1)$, we obtain conventional pure site (bond) percolation on the square lattice, and thus $p_{B,c}=p_{c}^{\mathrm{sq}}$ ($p_{A,c}=1/2$) along these lines, respectively (here, we used the fact that the percolation threshold $1/2$ for bond percolation on the square lattice has been exactly determined Kesten (1980)). This shows explicitly that due to the inequivalence of the two sublattices of the Lieb lattice, the phase diagram is not symmetric in the $(p_{A},p_{B})$-plane. In Ref. Tarasevich and Van der Mark (1999), the general case of mixed bond-site percolation is considered for various lattices, including the square lattice (mixed bond-site percolation on several lattices is also considered in Refs. Yanuka and Englman (1990); González et al. (2013); González-Flores et al. (2021); Torres et al. (2022)) . The phase diagram was numerically estimated, but only to comparably low accuracy, so that here we employed the bipartite Newman-Ziff algorithm in order to obtain more accurate values for the threshold line for sublattice-selective percolation on the Lieb lattice. The numerical results are provided in Table 2, and the corresponding phase diagram is shown in Fig. 9. Comparing the value of $p_{A,c}$ obtained for $p_{B}=1$ to the exact value ($1/2$) confirms the accuracy of our results. For estimating these numbers, $C$ was chosen identical to the square lattice value from the previous section, resulting in similarly fast convergence. Additionally, the previous estimate Oliveira et al. (2021) for the conventional ($p_{A}=p_{B}$) site percolation threshold $p_{c}^{\mathrm{Lieb}}$ on the Lieb lattice was improved using the Newman–Ziff algorithm. Along this line, we find that $C=0.68$ yields better convergence (however, even then the exponent $X\approx 1/\nu$ is significantly lower). The obtained value $p_{c}^{\mathrm{Lieb}}=0.7397120(17)$ is consistent with the previous best estimate Oliveira et al. (2021).
@@ V.1 Percolation on a chain ($z=2$)
In this case, percolation occurs only at the singular point $(p_{A},p_{B})=(1,1)$. To extract the critical exponents upon approaching the percolation point along different lines within the $(p_{A},p_{B})$-plane, we next examine appropriate state functions.
A cluster containing $s$ occupied sites with one empty site to either side is called an $s$–cluster. The number of $s$–clusters per lattice site $n_{s}$ is a useful state function. Another way of looking at $n_{s}$ is that it is the probability of an arbitrary site to be the left-end site of an $s$–cluster. This is the case if the site to the left is empty, then $s$ consecutive sites are occupied, followed by another empty site. For $p_{A}=p_{B}=p$, the state function would be $n_{s}=(1-p)p^{2}(1-p)$, but for $p_{A}\neq p_{B}$ one has to consider whether the left end site is from the $A$ or $B$ sublattice, and whether $s$ is even or odd. The probability that an arbitrary site is from the $A$ or $B$ sublattice is equal to $1/2$, and from there on it is a matter of counting the $A$– and $B$–sublattice sites in the cluster to find
@@
for even $s$, where $\bar{p}_{A}=1-p_{A}$, $\bar{p}_{B}=1-p_{B}$. The probability that an arbitrary site is part of an $s$–cluster is $n_{s}\cdot s$, so the probability that an arbitrary site is part of any cluster is given by
@@
which approaches 1 for $(p_{A},p_{B})\rightarrow(0,0)$ and diverges upon approaching $(1,1)$. We obtain the critical exponent $\gamma$ from analyzing the singular behavior of $S$ near $(1,1)$. Fixing $p_{B}=1$, the asymptotic singularity is obtained as
@@
To obtain the correlation length $\xi$, we first calculate the correlation function (pair connectivity function) $g(r)$, which equals the mean number of sites that are in the same cluster as a given occupied site, at a distance $r$ apart. Thus, there have to be $r-1$ consecutive occupied sites in between. For $p_{A}=p_{B}=p$ this is just $2p^{r}$ (2 for both directions), but for $p_{A}\neq p_{B}$ the parity of $r$ and the starting site have to be accounted for. The starting site is on the $A$ sublattice with probability $\frac{p_{A}}{p_{A}+p_{B}}$ and on the $B$ sublattice with probability $\frac{p_{B}}{p_{A}+p_{B}}$. Accounting for the parity of $r$ and the special case $r=0$, we obtain
$g(r)=\begin{dcases}1,&r=0,\\ \frac{4}{p_{A}+p_{B}}p_{A}^{(r+1)/2}p_{B}^{(r+1)/2},&\text{$r$ odd},\\ 2p_{A}^{r/2}p_{B}^{r/2},&\text{$r$ even}.\end{dcases}$ (17)
@@
The numerator in Eq. (18) can also be calculated, and for $p_{B}=1$ it scales as $(1-p_{A})^{-3}$ near the percolation threshold. Since $1/S$ scales as $(1-p_{A})$, the singularity of $\xi$ is thus obtained as
@@ V.2 Percolation threshold line for general $z$
The percolation condition on the Bethe lattice is that a chain of occupied sites reaches out to infinity. In other words, the system percolates if from some occupied site there exists a path of connected occupied sites that never ends (note that on the Bethe lattice such a path cannot form loops). If a chain of sites from a specific site up to some other site are occupied, at least 1 of the two neighbors going outwards also has to be occupied to have a percolating cluster. Hence, the mean number of occupied sites going outwards has to be at least one in order to form a percolating cluster. Furthermore, on the Bethe lattice the $A$ and $B$ sublattices alternate. The mean number of occupied sites going outwards can thus be calculated for $z$ neighbors by using two nested binomial distributions, one accounting for the mean number of $A$-sublattice sites and the other for the mean number of $B$-sublattice sites. This leads to the condition
@@
which can be simplified to yield the condition for the percolation transition,
@@
Note that this result is also correct for $z=2$, treated in the previous section. The percolation threshold line for $z=3$ is at $p_{A}\>p_{B}=1/4$, which is plotted in Fig. 10. The larger $z$ is, the bigger the area of the phase diagram that is percolating, but here only $z=3$ will be further examined.
@@ V.3 Scaling exponents for $z=3$
To obtain $P$, we first calculate the probability of a path not leading to infinity. Let $w_{A}$ ($w_{B}$) be the probability that an arbitrary $A$($B$) sublattice site does not lead to infinity, respectively. An $A$ sublattice site does not lead to infinity if the site itself is not occupied (given by the probability $1-p_{A}$) or if the two neighbors leading outwards do not lead to infinity (given by the probability $p_{A}w_{B}^{2}$), and similarly for a $B$ sublattice site. Altogether, this gives the recursive expressions
@@
This nonlinear system of equations has polynomials of degree four as solutions. The trivial solution is $w_{A}=w_{B}=1$, while the other real solution, denoted by $w^{P}_{A},w^{P}_{B}$ is rather lengthy, and given explicitly as a function of $p_{A}$ and $p_{B}$ in App. A. We note that for larger coordination number $z$, the corresponding recursion equations lead to higher-order equations for $w_{A}$ and $w_{B}$, for which closed expressions for the roots are known not not exist from Galois theory. For $p_{A}\>p_{B}<1/4$ there exists no infinite cluster, and thus the trivial solution applies, while for $p_{A}\>p_{B}\geq 1/4$ the other real solution is valid, i.e.,
$w_{A/B}=\begin{dcases}1,&4p_{A}\>p_{B}<1,\\ w^{P}_{A/B}(p_{A},p_{B}),&4p_{A}\>p_{B}\geq 1.\end{dcases}$ (25)
@@
since the probability that the center site is from the $A$ or $B$ sublattice is $1/2$ and the probability that the site is occupied and at least one neighbor leads to infinity equals $p_{A/B}(1-w_{B/A}^{3})$. The strength of the percolating cluster is plotted in Fig. 11.
@@
In Fig. 12, cuts of $P$ at three different values of $p_{B}$ are shown, along with the function $26p_{A}p_{B}-3/$ 2, which yields the leading asymptotic of $P$ for conventional site-percolation on the Bethe lattice Stauffer and Aharony (2018). We thus find that near the percolation threshold line, i.e., for $p_{A}\gtrsim 1/(4p_{B}),p_{B}\geq 1/4$, the scaling behavior
@@
Next, the mean cluster size $S$ will be calculated. For this, an arbitrary occupied starting site is considered and the mean mass of the cluster it belongs to is calculated. This approach is only valid below the percolation threshold line, i.e., for $14p_{A}p_{B}<$ 1, because otherwise the mass of the percolating cluster has to be accounted for as well. If the starting site belongs to the $A$($B$) sublattice, then let $S_{B}$ ($S_{A}$) be the contribution to the mean cluster size from each neighbor. The mean cluster size then is composed of the contribution from the center site plus the contribution from the three neighbors, so that
@@
The contributions $S_{A}$ and $S_{B}$ can be accessed recursively: If the starting site belongs to the $A$ sublattice, then a given neighbor only contributes a non–zero $S_{B}$ with probability $p_{B}$. The size of this contribution is the site itself plus $S_{A}$ for the two neighbors going outwards. With a similar reasoning for the $B$ sublattice, we thus obtain
@@
Both $S_{A}$ and $S_{B}$ diverge upon approaching the threshold condition $p_{A}p_{B}=\frac{1}{4}$. Hence, the behavior of $S$ near the percolation threshold line can be described by
@@
The correlation function $g(r)=g(\sqrt{r_{t}})$, for $14p_{A}p_{B}<$ 1, equals the mean number of occupied sites that are within the same cluster as a given occupied site, a topological distance $r_{t}$ apart – corresponding to the Euclidean distance $r=\sqrt{r_{t}}$. At $r_{t}=0$ the center site is already occupied, while at $r_{t}=1$ the site has 3 neighbors and from there on the number of neighbors doubles in each step, and the probabilities are otherwise the same as for the case of the chain, so that we obtain
@@ VI Conclusions
We presented an adapted Newman–Ziff algorithm to examine sublattice-selective percolation on bipartite lattices and we applied it to accurately determine the percolation threshold lines for the square lattice and the planar Lieb lattice. The latter case relates to mixed bond-site percolation on the square lattice, and our numerical results for the percolation threshold line refine previous estimates. Our numerical estimates for the critical exponents at the percolation transition are consistent with the universality class for percolation in $d=2$.
In addition, we examined sublattice-selective percolation on the Bethe lattice, which was in fact the geometry considered in the pioneering work on percolation by Flory in the context of polymerization Flory (1941). For this case, we obtained the exact percolation threshold line and critical exponents for $z=2$ and $z=3$, which are again in accord with the values for conventional percolation. Given that the sites of the square lattice have a uniform coordination number of $z=4$, it may be interesting to compare its phase diagram to that of the Bethe lattice with $z=4$. Such a comparison is shown in Fig. 13, illustrating the substantially larger percolating regime for the latter case. Indeed, the absence of closed loops for a chain of connected sites leads to an overall prior appearance of the percolating cluster upon increasing the site occupation probabilities on the Bethe lattice. We checked explicitly, that a rescaling of the analytic expression for the percolation threshold line for the $z=4$ Bethe lattice to match the ends points at $p_{A}=1$ and $p_{B}=1$ does not fit the full numerical data.
@@
Related to the last point, we note that the authors of Ref. Tarasevich and Van der Mark (1999) considered various proposed analytic expressions for the general functional form of the percolation threshold line for mixed bond-site percolation. They observed clear systematic deviations from the numerical data for low-coordinated lattices. Based on our refined data on the Lieb lattice, we can also exclude the validity of these expressions for the square lattice (cf. Appendix B). It thus remains an interesting open question for further research to derive analytical expressions for the percolation threshold line for sublattice-selective as well as mixed bond-site percolation, respectively. We hope that our refined numerical data will be valuable as a benchmark for such investigations. Finally, let us note that sublattice-selective percolation can also be related to the problem of antisite defect percolation, such as that considered in Ref. Tarasevich and Manzhosova (2003) for the case of a simple cubic lattice. Certainly, the adapted Newman-Ziff algorithm can be of use also for further investigations of such generalized percolation problems.